OpenStudy (beccab003):

Simplifying radicals help! Medal to best answer. question will be posted below.

OpenStudy (beccab003):

\[(4-\sqrt{3})^{2}\]

OpenStudy (beccab003):

my answer: \[19-8\sqrt{3}\]

OpenStudy (beccab003):

@DanJS

OpenStudy (danjs):

\[(4-\sqrt{3})*(4-\sqrt{3}) = \]

OpenStudy (beccab003):

I used foil for that.

OpenStudy (danjs):

\[4*4 - 4\sqrt{3}-4\sqrt{3}+\sqrt{3}\sqrt{3}\]

OpenStudy (danjs):

\[16 - 8\sqrt{3} + 3 = 19-8\sqrt{3}\]

OpenStudy (danjs):

good work

OpenStudy (beccab003):

Thanks! I have another that i have gone through multiple times but can't get it right... do you want me to post in a new thread?

OpenStudy (danjs):

sure